# Global-Local Progressive Integration Network for Blind Image Quality Assessment

### v1

@@ Abstract
Vision transformers (ViTs) excel in computer vision for modeling long-term dependencies, yet face two key challenges for image quality assessment (IQA): discarding fine details during patch embedding, and requiring extensive training data due to lack of inductive biases. In this study, we propose a Global-Local progressive INTegration network for IQA, called GlintIQA, to address these issues through three key components: 1) Hybrid feature extraction combines ViT-based global feature extractor (VGFE) and convolutional neural networks (CNNs)-based local feature extractor (CLFE) to capture global coarse-grained features and local fine-grained features, respectively. The incorporation of CNNs mitigates the patch-level information loss and inductive bias constraints inherent to ViT architectures. 2) Progressive feature integration leverages diverse kernel sizes in embedding to spatially align coarse- and fine-grained features, and progressively aggregate these features by interactively stacking channel-wise attention and spatial enhancement modules to build effective quality-aware representations. 3) Content similarity-based labeling approach is proposed that automatically assigns quality labels to images with diverse content based on subjective quality scores. This addresses the scarcity of labeled training data in synthetic datasets and bolsters model generalization. The experimental results demonstrate the efficacy of our approach, yielding 5.04% average SROCC gains on cross-authentic dataset evaluations. Moreover, our model and its counterpart pre-trained on the proposed dataset respectively exhibited 5.40% and 13.23% improvements on across-synthetic datasets evaluation [^1] The codes and proposed dataset will be released at here..
@@
Blind image quality assessment, dataset construction, vision transformer, features fusion.
@@ I Introduction
Objective image quality assessment (IQA) aims to enable automated evaluation of visual image quality in alignment with human perception. IQA plays a crucial role in various image processing pipelines, such as restoration, medical imaging and retrieval, ensuring optimal visual experiences across diverse applications. IQA methods can be classified into three groups based on the availability of reference information: full-reference (FR-IQA), reduced-reference (RR-IQA), and no-reference IQA (NR-IQA) or blind IQA (BIQA) [1]. Owing to the unavailability of reference information in real-world scenarios, BIQA methods have attracted considerable attention and research efforts in the IQA domain.
Traditional BIQA methods [2, 3, 4, 5, 6] rely on hand-crafted features, such as discrete wavelet coefficients [2] and locally normalized luminance coefficients [6], for training using traditional learning algorithms, such as support vector regression (SVR) [3] and multivariate Gaussian models (MVG) [5]. However, these methods are limited by the need for extensive domain expertise and the inability to fully capture complex distortions in real-world images, affecting accuracy and generalization. In contrast, deep learning-based BIQA methods [7, 8, 9, 10, 11, 12, 13] leverage neural networks to learn distortion features and image quality end-to-end, demonstrating promising performance and becoming a highly sought-after research direction.
The effectiveness of deep learning-based BIQA methods depends on two critical factors: model and data. Convolutional neural networks (CNNs) are prevalent and effective tools for feature learning and representation. In the process, convolutional kernels slide across the spatial dimensions of the input images, focusing on local windows to detect visual patterns and transform the data into discriminative feature spaces. In developing CNNs-based BIQA models, quality-aware features can be learned through various distortion-related optimization paradigms, such as distortion type classification [9] and distortion strength ranking [14]. These features are then inputted into regression modules like multi-layer perceptrons (MLPs) to predict image quality scores. To improve the generalization of BIQA models trained on limited IQA databases, researchers have introduced augmentation techniques, such as patch cropping to enlarge the training set and synthesizing distorted images with varied distortion types and parameters [10].
Recently, vision Transformers (ViTs) have garnered substantial interest across diverse vision applications. Researchers have also investigated their potential in learning image quality. Generally, ViTs-based FR-IQA models [15] can achieve strong performance, benefitting from access to reference information. However, there are still two limitations associated with ViTs-based BIQA models. First, some studies leverage ViTs as regression modules to learn image quality from CNN-extracted features [16, 17]. However, such approaches inherently limit the representational capability of the CNN-extracted features that fail to fully encode long-range dependencies. Sencod, directly utilizing ViTs [18, 19, 20] as feature extraction modules may also pose challenges due to the the patch embedding process, which discretizes the input image into visual tokens, where each token represents a low-resolution image patch rather than an individual pixel, leading to a loss of local fine-grained spatial details.
Moreover, despite ViTs demonstrating proficiency in capturing global dependencies, their lack of inductive biases necessitates larger datasets to achieve comparable performance [21]. However, the synthetic distortion datasets, such as [22, 23, 24], commonly employed for evaluating BIQA models often suffer from limited content diversity. For instance, TID2013 [23] comprises a mere 3,000 images across 25 content types, while KADID-10k [24] offers 10,125 images with 81 content types. The development of extensive and varied IQA datasets is constrained by the high costs and complexities of subjective quality assessments, limiting ViTs’ efficacy in understanding diverse image distortions. While prior works have explored data augmentation, such as using injected distortion levels or FR-IQA algorithms for quality annotation, the methods only deliver coarse information or exhibit distortion type biases [25]. Thus, it is crucial to devise more accurate labeling techniques to improve the robustness of BIQA models.
This paper addresses two critical aspects of IQA: model architecture and training data. Regarding model design, we propose a Global-Local progressive INTegration network, termed GlintIQA, that combines ViT and CNNs-based extractors to capture multi-grained feature representations, which are subsequently fused through channel-wise attention and spatial enhancement modules. The progressive integration approach efficiently handles diverse multi-grained features simultaneously, with each integration stage involving spatial feature enhancement and channel importance optimization to refine the resulting fused representations. Regarding training data, we exploit the correlation between content similarity and quality scores to facilitate automated labeling, which directly utilizes subjective scores to better capture the distribution of human quality perceptions. The main contributions of this paper are threefold:
  • We proposed GlintIQA that employs a hybrid feature extraction scheme combining a ViT-based global feature extractor (VGFE) and a CNN-based local feature extractor (CLFE). This synergistic architecture effectively captures both global coarse-grained and local fine-grained features, addressing the patch-level information loss and inductive bias issues inherent in ViTs.
  • We proposed an efficient feature integration strategy to handle multi-grained representations from the hybrid extractors. The strategy utilizes multiple kernel embeddings for spatial alignment, and progressively integrates fine and coarse-grained features using channel-wise attention and spatial enhancement modules.
  • We proposed a content similarity-guided labeling approach that leverages the correlation between content similarity and quality scores for automated annotation of training data. This method directly incorporates subjective scores, better reflecting human perception distribution and improving annotation accuracy over techniques based on FR-IQA predictions.
The paper is structured as follows: Section II reviews BIQA methods regarding models and data. Section III details the proposed global-local integration network. Section IV outlines dataset construction procedures, including image degradation and label assignment. Section V presents extensive comparisons with existing BIQA methods across ten IQA datasets and ablation studies validating the effectiveness. Finally, Section VI provides a conclusion.
II Related Work
In this section, related work on deep learning-based BIQA methods has been reviewed from both the model architecture and training data perspectives.
II-A Architecture of BIQA Models
@@
Over the past decade, researchers have developed diverse neural networks, such as CNNs, and ViTs, as well as various architectures, such as single-stream, dual-stream, and duo-stream structures, to learn image distortion representations and regress image quality scores. CNNs have emerged as the dominant architecture for developing diverse BIQA models. Early BIQA models [7, 8, 11] primarily employed single-stream architectures, stacking CNN layers and linear layers to regress image quality scores in an end-to-end manner. To enhance assessment accuracy, researchers have designed dual-stream architectures that enable more effective feature extraction and optimized quality representation learning. For instance, Wang et al. [26] integrated distorted images with their corresponding gradient maps to augment the extraction of structural detail features. Ma et al. [9] introduced an auxiliary distortion recognition branch into the quality prediction network, accounting for the impact of diverse distortion types on image quality perception. Zhang et al. [12] presented a dual-stream architecture that integrates semantic and distortion perception, enabling adaptation to both authentic and synthetic distortion assessment scenarios. On the other hand, Liu et al. [10] employed Siamese networks to establish pairwise rankings among images exhibiting varying levels of distortion for learning the relative quality representations. To enrich feature extraction, Wu et al. [27] proposed a multi-stream structure that simultaneously extracts features from distorted images, reconstructed images, error maps, and structural degradation maps. Furthermore, Wu et al. [28] adopted hierarchical structures that employ modules across different stages of the model to capture features across diverse scales.
@@
CNN-based BIQA models have shown significant progress, with the transition from single-stream to multi-stream and hierarchical structures demonstrating a clear trend towards richer and more robust feature extraction techniques. However, despite these advancements, CNNs’ inherent local processing nature [29] can limit their capacity to model global dependencies within images. With the success of transformer models in visual tasks, several studies [17, 16, 30, 31, 32, 33, 34] have explored their application in BIQA. For instance, You et al. [17] leveraged Transformer to establish a mapping from semantic features to quality scores. Golestaneh et al. [16] further constrained the features to represent quality consistency under flipping and learned the relative quality ordering among images. Zhu et al. [30] introduced a saliency map within the feature space to guide the Transformer’s attention towards regions of interest during the regression phase. These works have been focused on developing CNNs-based feature extractors and Transformer-based regressors. In addition, studies utilizing Transformers as feature extraction backbones has demonstrated promising results [33, 34]. Wang et al. [33] built a swin Transformer-based siamese network trained jointly on regression and pairwise ranking tasks. Qin et al. [34] proposed an attention mechanism on transformer-encoded features, mimicking diverse human evaluations for quality prediction from various perspectives. While ViTs-based backbones excel at capturing global features, the patch embedding process makes them insensitive to intra-patch details and lacks the inductive biases inherent in CNN architectures. Therefore, the development of hybrid models combining CNNs and ViTs has the potential to leverage the advantages of both architectures, offering a comprehensive feature extraction approach that addresses both local and global aspects of image quality. This hybrid strategy could mitigate the inductive biases of ViTs, resulting in more accurate and robust BIQA models.

### v2

# No-Reference Image Quality Assessment with Global-Local Progressive Integration and Semantic-Aligned Quality Transfer

@@ Abstract
Accurate measurement of image quality without reference signals remains a fundamental challenge in low-level visual perception applications. In this paper, we propose a global-local progressive integration model that addresses this challenge through three key contributions: 1) We develop a dual-measurement framework that combines vision Transformer (ViT)-based global feature extractor and convolutional neural networks (CNNs)-based local feature extractor to comprehensively capture and quantify image distortion characteristics at different granularities. 2) We propose a progressive feature integration scheme that utilizes multi-scale kernel configurations to align global and local features, and progressively aggregates them via an interactive stack of channel-wise self-attention and spatial interaction modules for multi-grained quality-aware representations. 3) We introduce a semantic-aligned quality transfer method that extends the training data by automatically labeling the quality scores of diverse image content with subjective opinion scores. Experimental results demonstrate that our model yields 5.04% and 5.40% improvements in Spearman’s rank-order correlation coefficient (SROCC) for cross-authentic and cross-synthetic dataset generalization tests, respectively. Furthermore, the proposed semantic-aligned quality transfer further yields 2.26% and 13.23% performance gains in evaluations on single-synthetic and cross-synthetic datasets [^1] The codes and proposed dataset will be released at here..
@@
Image quality assessment, dataset construction, label transfer, feature integration, vision transformer.
@@ I Introduction
As a crucial measurement component in modern imaging systems, objective image quality assessment (IQA) enables quality control and optimization across various applications including medical imaging, visual inspection, and multimedia services. IQA aims to enable automated quantitative evaluation of visual image quality in alignment with human perception. Based on the availability of reference signals, IQA methods are categorized into three paradigms: full-reference (FR-IQA), reduced-reference (RR-IQA), and no-reference IQA (NR-IQA) [1]. Among these, NR-IQA has gained significant research attention due to its practical applicability in real-world scenarios where reference images are unavailable.
Traditional NR-IQA methods [2, 3, 4, 5, 6] extract predefined measurement features, e.g., wavelet coefficients [2] and normalized luminance statistics [5], which are then calibrated using regression models like support vector regression (SVR) [3] and multivariate Gaussian models (MVG) [4]. However, these methods face limitations in measurement accuracy and reliability due to their dependence on manually designed features and inability to comprehensively characterize complex image degradations. In contrast, deep learning-based NR-IQA models [7, 8, 9, 10, 11, 12, 13, 14] demonstrate superior performance by end-to-end learning quality-relevant features from training data.
@@
The effectiveness of deep learning-based NR-IQA methods depends on two main components: measurement model and training data. Convolutional neural networks (CNNs) serve as effective tools for feature extraction and quality assessment. Through spatially-sliding convolutional kernels, the networks analyze local image regions to detect visual patterns and transform them into discriminative quality-relevant features. Early CNNs-based NR-IQA models[10, 9] primarily employed single-stream architectures, stacking CNN layers and linear layers to regress image quality scores in an end-to-end manner. To enhance the assessment accuracy, researchers have developed increasingly sophisticated architectures [15, 7, 12, 16, 8, 17, 18]. Wang et al.[15] integrated distorted images with their corresponding gradient maps to augment the extraction of structural detail features. Ma et al.[7] introduced an auxiliary distortion recognition branch into the quality prediction network, accounting for the impact of diverse distortion types on image quality perception. Zhang et al.[12] and Pan et al.[16] presented a dual-stream architecture that integrates semantic and distortion perception, enabling adaptation to both authentic and synthetic distortion measurement scenarios. Liu et al.[8] employed Siamese networks to establish pairwise rankings among images exhibiting varying levels of distortion for learning relative quality representations. Further advancing architectural complexity, Wu et al.[17] proposed a multi-stream structure that simultaneously extracts features from distorted images, reconstructed images, error maps, and structural degradation maps. Wu et al.[18] adopted hierarchical structures that employ modules across different stages of the model to capture features across diverse scales.
@@
However, despite these advancements, CNNs’ inherent local processing nature[19] can limit their capacity to model global dependencies within images. This limitation has driven researchers to explore the use of vision Transformers (ViTs) that have shown exceptional ability in capturing long-range spatial correlations. The integration of ViTs into NR-IQA has followed two distinct approaches. The first approach uses ViTs primarily as regression modules while retaining CNNs for feature extraction[20, 21, 22]. You et al.[20] leveraged Transformer to establish a mapping from semantic features to quality scores. Golestaneh et al.[21] further constrained the features to represent quality consistency under flipping and learned the relative quality ordering among images. Zhu et al.[22] introduced a saliency map within the feature space to guide the Transformer’s attention towards regions of interest during regression. The second approach employs ViTs as the primary feature extraction backbone[23, 24]. Wang et al.[23] built a swin Transformer-based siamese network trained jointly on regression and pairwise ranking tasks. Qin et al.[24] proposed an attention mechanism on transformer-encoded features, mimicking diverse human evaluations for quality prediction from various perspectives. However, both approaches face challenges in NR-IQA applications. The first approach constrains feature representation by limiting ViTs to post-processing CNNs-extracted features [21, 20]. The second approach encounters difficulties due to ViTs’ patch-based tokenization [25, 26], where each token represents a low-resolution image patch, potentially compromises the measurement of spatial details critical for quality evaluation.
Moreover, while ViTs excel at capturing global dependencies, their lack of inductive biases necessitates larger datasets to achieve comparable performance [27]. However, commonly used synthetic datasets [28, 29, 30] often lack content diversity. For instance, TID2013 [29] contains only 3,000 images across 25 content types, and KADID-10k [30] provides 10,125 images with 81 content types. Expanding these datasets is constrained by the high costs and complexities of subjective quality evaluation procedures, which impacts ViTs’ effectiveness in understanding diverse image degradations. Researchers have attempted to address the data limitations through various augmentation techniques. The patch-based approach randomly selects small patches (e.g., 112$\times$112 as used in [10]) from the full images. To mitigate potential label bias from small patches, Kim et al.[31] utilized patch-level quality scores from FR-IQA methods as intermediate regression targets. Several studies[7, 8, 17, 18] have created large-scale datasets for pre-training NR-IQA models. Ma et al.[7] synthesized images with different distortions using artificial degradation for training distortion classification networks. Liu et al.[8] leveraged injected distortion levels of synthetic images as quality rankings. Previous attempts at data augmentation, including distortion level injection[8] or FR-IQA algorithm annotation[18], only provide coarse quality indicators or show distortion-specific biases[32]. While FR-IQA algorithms can generate fine-grained image quality scores, they often exhibit biases towards certain distortion types and rely primarily on low-level feature discrepancies, overlooking the crucial role of high-level semantic content in visual quality perception.
Based on the issues stated above, this paper focuses on two critical aspects: model architecture and training data. Our main contributions are three-fold:
  • To overcome architectural limitations, we propose a Global-Local progressive Integration model (termed GlintIQA) for image quality assessment. This architecture integrates ViT-based global feature extraction (VGFE) with CNN-based local feature extraction (CLFE), leveraging ViTs’ strength in capturing long-range dependencies through self-attention mechanisms while preserving CNNs’ capability in quantifying fine-grained spatial details. This dual-stream design effectively overcomes both the patch-level information loss and inductive bias constraints in pure ViT-based approaches and the local context constraints in CNN-only architectures.
  • To facilitate effective feature integration, we develop an efficient progressive integration strategy that considers both channel-wise and spatial information across features of different granularities. Our approach introduces cascaded channel-wise self-attention (CWSA) modules that dynamically adjust feature importance across different granularities, coupled with spatial interaction enhancement modules (SIEM) that strengthen the interactions between cross-scale local and global features, leading to improved quality assessment accuracy.
  • To address the challenges of data scarcity and annotation limitations in model training, we propose a semantic-aligned quality transfer (SAQT) method for automated image quality annotation. This approach exploits perceptual correlations between semantically similar images under identical distortion conditions to transfer subjective quality scores. Through statistical validation, we demonstrate the efficacy of the SAQT-based annotation method and establish the SAQT-IQA dataset, which yields substantial improvements in model generalization capabilities when employed for training.
The paper is structured as follows: Section II describes the quality labeling method and the construction of the proposed SAQT-IQA dataset. Section III details the proposed GlintIQA model. Section IV presents comprehensive experimental results and ablation studies across multiple IQA benchmark databases. Finally, Section V draws a conclusion.
(a) CSIQ [33]
(b) TID2013 [29]
(c) KADID-10k [30]
(d) LIVE-MD [34]
Figure 1: The correlation between image semantic distance and quality scores is demonstrated under identical distortion conditions across four IQA datasets.
II The proposed semantic-aligned quality transfer method
We propose semantic-aligned quality transfer method that leverages the perceptual correlation of similar content under identical distortions [35] for labeling image quality score. We begin with a statistical analysis to validate the feasibility of our method (section II-A). Then, we detail the construction of a large-scale synthetic dataset with this method (section II-B).
II-A Statistcal Analysis
To validate the relationship between semantic similarity and quality perception, we conducted a comprehensive statistical analysis across four IQA benchmark datasets. These datasets consist of pristine images and their corresponding distorted versions, each associated with mean opinion scores (MOS). Fig. 1 illustrates the relationship between semantic distance (on the x-axis) and the correlation of quality scores (on the y-axis) across these datasets. Specifically, the x-axis represents the cosine distance between semantic feature vectors (extracted by ResNet50) of pristine image pairs, while the y-axis indicates the correlation of quality scores for the distorted images corresponding to these pristine pairs. The results reveal that for image pairs with larger semantic distances, the correlation between their quality scores exhibits greater variability. In contrast, as the semantic distance between images decreases, the correlations become increased and more consistent, approaching a value of 1 as the distance nears zero. This trend is consistently observed across all four datasets, which reinforces the conclusion that images with similar content tend to exhibit comparable visual quality labels when subjected to identical distortion conditions. Building on these insights, the proposed labeling method leverages human-annotated images to transfer quality labels to degraded images with similar content, enabling the construction of a large-scale synthetic dataset guided by semantic similarity. Next, we describe the specific steps for constructing the expanded dataset.
II-B Dataset Construction
II-B1 Preliminary Data Preparation
To create the dataset, we meticulously selected 50,000 high-quality images from the KADIS-700k [30] and applied 25 types of distortions with five intensity levels to each image. The distortion types were selected to be consistent with those present in KADID-10k [30]. Finally, this process produced 6,250,000 degraded images. The quality of the generated images is annotated using MOS derived from the KADID-10k IQA database. Let $\mathcal{X}=\left\{(\textbf{X}_{i,d},\textbf{M}_{i,d})\right\}_{i=1}^{P}$ denote the set of pristine images and their corresponding degraded versions, where $\textbf{X}_{i,d}=\left\{x^{t,l}_{i,d}|t\in\mathcal{T},l\in\mathcal{L}\right\}$ represents the degraded images of the $i$-th pristine image, and $\textbf{M}_{i,d}=\left\{m^{t,l}_{i,d}|t\in\mathcal{T},l\in\mathcal{L}\right\}$ are their corresponding MOS values, scaled to a range between 0 and 9. The high-quality images are represented by the set $\textbf{Y}=\left\{Y_{j}\right\}_{j=1}^{H}$, where $Y_{j}$ refers to the $j$-th high-quality image and $H$ is the total number of high-quality images.
Figure 2: Semantic distance analysis and dataset validation. Upper: PLCC distribution across semantic distances for KADID-10k (mean and standard deviation). Lower: Distribution of semantic distances in the proposed SAQT-IQA dataset.
Figure 3: The procedure of dataset construction based on the proposed semantic-aligned quality transfer method.
II-B2 Semantic-Aligned Quality Transfer (SAQT)-IQA dataset
The process of constructing the dataset is illustrated in Fig. 3, which involves the following three steps: Each high-quality image $Y_{j}$ and candidate pristine image $X_{i}$ are mapped into a high-dimensional semantic space through a feature extraction function $\mathcal{F}$, which is based on a pre-trained ResNet101 [36], excluding the final linear layer. The extracted feature vectors $\mathcal{F}(X_{i})$ and $\mathcal{F}(Y_{j})$ belong to the space $\mathbb{R}^{2048}$. These vectors capture high-level semantic information for the measurement of content similarity between images. (b) Content similarity determination: To quantify the semantic similarity between $Y_{j}$ and the set of candidate pristine images $\{X_{i}\}_{i=1}^{P}$, we define a semantic distance function $d_{s}(\cdot)$:
$d_{s}(\mathcal{F}(X_{i}),\mathcal{F}(Y_{j}))=1-\frac{\langle\mathcal{F}(X_{i})\cdot\mathcal{F}(Y_{j})\rangle}{||\mathcal{F}(X_{i})||_{2}\cdot||\mathcal{F}(Y_{j})||_{2}},$ (1)
[…]
$X_{\hat{i}}=\arg\min\limits_{\footnotesize X_{i}\in\{X_{i}\}_{i=1}^{P}}d_{s}(\mathcal{F}(Y_{j}),\mathcal{F}(X_{i})),$ (2)
This step identifies the pristine image $X_{\hat{i}}$ that most closely matches the content of $Y_{j}$ in terms of semantic features. (c) Quality transfer operator: Once the most semantically similar pristine image $X_{\hat{i}}$ is identified, the next step involves transferring the quality labels from the pristine image to the high-quality image $Y_{j}$. This is formalized using the quality transfer function $\mathcal{T}$, which maps the high-quality image $Y_{j}$ to its corresponding degraded version $y^{t,l}_{j,d}$ and assigns the appropriate quality label $s^{t,l}_{j,d}$:
$\mathcal{T}:(Y_{j},X_{\hat{i}})\mapsto\left\{\begin{aligned} &\text{Degradation:}\quad y^{t,l}_{j,d}=\mathcal{D}_{t,l}^{\hat{i}}(Y_{j})\\ &\text{Quality Label:}\quad s^{t,l}_{j,d}=m^{t,l}_{\hat{i},d}\end{aligned}\right.$ (3)
Here, $\mathcal{D}_{t,l}^{\hat{i}}(Y_{j})$ denotes the degradation operation applied to $Y_{j}$, producing a distorted version $y^{t,l}_{j,d}$. The MOS value $m^{t,l}_{\hat{i},d}$, associated with the degraded version $x^{t,l}_{\hat{i},d}$ of $X_{\hat{i}}$, is then transferred to annotate the degraded image $y^{t,l}_{j,d}$. Finally, the generated dataset can be represented as $\mathcal{Y}=\left\{(\textbf{Y}_{j,d},\textbf{S}_{j,d})\right\}_{j=1}^{H}$, where $\textbf{Y}_{j,d}=\left\{y^{t,l}_{j,d}|t\in\mathcal{T},l\in\mathcal{L}\right\}$ and $\textbf{S}_{j,d}=\left\{s^{t,l}_{j,d}|t\in\mathcal{T},l\in\mathcal{L}\right\}$.
The effectiveness of the SAQT-IQA dataset is further validated through an analysis of semantic distances and their correlation with MOS values. Specifically, the relationship between semantic distances and PLCC values is examined by partitioning the distance range into 0.07 intervals and calculating the mean and standard deviation within each interval based on the results shown in Fig. 1(c). In Fig.  2, the upper plot illustrates this correlation of KADID-10k, where a gradual decline in correlation is observed as semantic distance increases. Importantly, the majority of PLCC values remain elevated ($>$0.9) for semantic distances below ($<$0.4). The lower plot presents the semantic distance distribution of our constructed dataset, exhibiting an normal distribution ($\mu$=0.206, $\sigma$=0.055). This distribution serves two purposes: image pairs with distances below the mean $\mu$ promote label accuracy due to semantic proximity, while those above the mean $\mu$ enhance content diversity. Crucially, PLCC values stay consistently high ($>$0.9) across these distances, confirming the reliability of our label transfer strategy and validating the effectiveness of the constructed dataset. The proposed SAQT method generates quality scores that closely align with subjective human annotations, considering the influence of image content on quality perception. In contrast, the FR-IQA approach largely neglects high-level semantics and disregards content-related perceptual factors in quality measurement.